OpenStudy (anonymous):

Could someone please help me? Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 4 h, and Car B traveled the distance in 4.5 h. Car B traveled 5 mph slower than Car A. How fast did Car B travel?

OpenStudy (anonymous):

40mph :)

OpenStudy (anonymous):

Could you please tell me how to get the answer?

OpenStudy (anonymous):

4r = 4.5(r-5) 22.5/.5 = r = 45mph slower car traveled at 40 mph. 45 - 5 = 40 mph